Savannah has $9,900. She invested her savings in two banks. One bank is paying 5% simple interest and the other 6%. How much money did Savannah invest in each bank if the annual interest is $570.00?
Answer by jorel1380(3719)

p+q=9900
.05p+.06q=570
5p+6q=57000
5p+5q=49500
q=7500
p=2400
Savannah has $2400 invested at 5% and $7500 in at 6%..